OpenStudy (anonymous):

What is the mean of the following data without the outlier? 20, 25, 30, 3

OpenStudy (anonymous):

19.5 25 18.75 20

OpenStudy (anonymous):

this time you dont have to sort :)

OpenStudy (anonymous):

ya hahahaha

OpenStudy (anonymous):

add up all the numbers in the set and divide by how many members are in the set :) can you do that?

OpenStudy (anonymous):

ya thx

OpenStudy (anonymous):

(20+ 25+ 30+ 3 )/4

OpenStudy (anonymous):

so its 18.75?

OpenStudy (anonymous):

19.5 :(

OpenStudy (anonymous):

oh :(

OpenStudy (anonymous):

(20+ 25+ 30+ 3 )/4= 78/4= do parenthesis first 19.5 divide by 4